A flask contains 1.2 moles of CH4 at 25C and 74.6 mm pressure.


If 4 gms of sme gaas is sent into the flask, find its pressure.

Grade:

1 Answers

Chetan Mandayam Nayakar
11 years ago

4 gms=0.25 moles

P1/P2 =n1/n2 =1.45/1.2

answer=(1.45/1.2)*74.6 mm